hi i need help with solving graphing linear equation with two variables i do not know how to plot the correct number to graph for x the question is -5x+10y=-50
Answer by jim_thompson5910(35256)

When x = 0, y is...

-5x+10y = -50
-5*0+10y = -50
0+10y = -50
10y = -50
y = -50/10
y = -5


So the y-intercept is (0,-5)

-------------------------------------------------------

When y = 0, x is...

-5x+10y = -50
-5x+10*0 = -50
-5x+0 = -50
-5x = -50
x = -50/(-5)
x = 10


So the x-intercept is (10,0)

-------------------------------------------------------

So to graph -5x+10y=-50, you plot the two points (0,-5) and (10,0)

Then you draw a straight line through the two points. This line will extend infinitely in both directions.